OpenStudy (anonymous):

Verify the identity. ((cos x)/1+sin x) + ((1+sinx)/cos x) =2secx

OpenStudy (anonymous):

\[\frac{ cosx }{ 1+\sin x }+\frac{ 1+\sin x }{ \cos x }=2secx\]

OpenStudy (anonymous):

I don't know if I'm doing this right. I used LCD.

OpenStudy (anonymous):

\[LEFT SIDE: \frac{ \cos^2x +(1+\sin x)(1+\sin x)}{ \cos(1+sinx) }\]

OpenStudy (anonymous):

\[ \frac{ \cos^2x +1+2sinx+sin^2x}{ \cos(1+sinx) }\]

OpenStudy (mertsj):

\[\frac{(\cos x)(1-\sin x)}{(1+\sin x)(1-\sin x)}+\frac{(1+\sin x)(\cos x)}{\cos ^2x}=\frac{\cos x-\cos x \sin x}{1-\sin ^2x}+\frac{\cos x+\cos x \sin x}{\cos ^2x}\]

Nnesha (nnesha):

change cos^2 by 1-sin^2x identity

OpenStudy (anonymous):

Wait, am I doing it right?

OpenStudy (mertsj):

\[\frac{2\cos x}{\cos ^2x}=\frac{2}{\cos x}=2\sec x\]

OpenStudy (anonymous):

@Mertsj What did you do first?

OpenStudy (mertsj):

Multiply first fraction by (1-sinx)/(1-sinx)

OpenStudy (mertsj):

Multiply second fraction by cosx/cosx

OpenStudy (anonymous):

Ohhh is it like multiplying the conjugate?

OpenStudy (mertsj):

it's really setting up cos^2x in both denominators

OpenStudy (mertsj):

Because 1-sin^2x=cos^2x

Nnesha (nnesha):

\[\huge\rm \frac{ \color{Red}{\cos^2x} +1 +2sinx + \sin^2 }{ cosx(1+sinx) }\] \[\large\rm \frac{ \color{Red}{1-sin^2x} +1 +2sinx + \sin^2 }{ cosx(1+sinx) }\] sin^2 cancel each other out you will get \[\large\rm \frac{ \color{Red}{1\cancel{-sin^2x}} +1 +2sinx +\cancel{ \sin^2} }{ cosx(1+sinx) }\]

OpenStudy (anonymous):

Ohhh so what I was doing is wrong. Haha! Thanks!

Nnesha (nnesha):

i don't think so..

OpenStudy (mertsj):

Not wrong, such not very effective.l

OpenStudy (anonymous):

Like it the solution would be longer? :D

Nnesha (nnesha):

\[\huge\rm \frac{ 2+2sinx }{ cosx(1+sinx) }\] take out 2 \[\rm \frac{ 2(1+sinx) }{ cosx(1+sinx)}\]

OpenStudy (anonymous):

There's no "it" in my last sentence. Haha.

OpenStudy (mertsj):

Typically there are a variety of ways to solve these identities.

OpenStudy (anonymous):

Ohhhh @Nnesha thank you so much! :) You did the whole solution for me.

Nnesha (nnesha):

true but i think the way you were doing is easy :D

OpenStudy (anonymous):

That's why this is harder because there are so many ways to do it and it's so frustrating haha!

Nnesha (nnesha):

and like mertsj said there are more than 2 ways to verify the identities